OpenStudy (anonymous):

giving madals How much interest is earned on $2,700 in one year if it is invested at 3%? $8.10 $81.00 $810.00

OpenStudy (solomonzelman):

Simple, or compound some number of times per year ?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

stop giving direct ans @zDior_VII

OpenStudy (yanasidlinskiy):

Use the formula for compound interest:) \(\huge{F = 2700(1+0.03)^1}\) \(\huge{F = 2781}\) Which rounds too..

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

subtract those two values and u will get the interest which is 81

OpenStudy (solomonzelman):

yana nice try, but \(\LARGE\color{blue}{ \bf Ft=P(1+\frac{percent~~~rate}{n})^{nt} }\)

OpenStudy (solomonzelman):

who told you that it is compound once per year ?

OpenStudy (anonymous):

so the seconde one

OpenStudy (yanasidlinskiy):

So interest = \(\huge2781-2700 =\) ? Yes. It's 81. Am I wrong? Really?

OpenStudy (solomonzelman):

(If it is compound at all)

OpenStudy (yanasidlinskiy):

I had the same exact not long ago.

OpenStudy (yanasidlinskiy):

*question*

OpenStudy (solomonzelman):

81 is simple interest, or compound once per year. For 1 year, those 2 are same.

OpenStudy (anonymous):

i wish i could give more than one medal

OpenStudy (anonymous):

at the end it is still 81

OpenStudy (yanasidlinskiy):

ok, so are they going to be different for next years then? I don't think soo..

OpenStudy (anonymous):

@zDior_VII dont force us to report u!

OpenStudy (solomonzelman):

She didn't know:)

OpenStudy (anonymous):

explain if u know the ans pls

OpenStudy (solomonzelman):

Or she did... who knows :P

OpenStudy (anonymous):

im a dude

OpenStudy (yanasidlinskiy):

I do know that they know:) I have already told them in the other simple/easy question:)

OpenStudy (solomonzelman):

I more often see questions that compare interests. Compound interest is view as an exponential function and simple interest is view as a linear function.

OpenStudy (solomonzelman):

Well if they know that they know (sound weird), you can report if you want....

OpenStudy (anonymous):

wtvs VII

OpenStudy (yanasidlinskiy):

Hahha:) Solomon:) "Well if they know that they know (sounds weird)" I just added the s, to make it look different lol:)

OpenStudy (solomonzelman):

yeah I left out the s... :)

OpenStudy (yanasidlinskiy):

Hahahha:) Yea.I caught ya on that one:)!!!!

OpenStudy (solomonzelman):

Time for compound interest, withdrawing from the bank account a certain (same) amount per year, or something that is a bit harder ? or not :) (?)

OpenStudy (yanasidlinskiy):

That doesn't really make sense to me..\(\huge{\uparrow}\)

OpenStudy (solomonzelman):

Good luck with interests:) I would teach you something if you want. I taught my sister everything about Arithmetic and geometric sequence in 5 minutes :PPPP \(\normalsize\color{blue}{ \rm {Let's~~~✂~~the~~above~~comment~~ } \Huge{☻ } }\)

OpenStudy (yanasidlinskiy):

Hahha:) Go ahead and teach me:) I would rather have someone teach me that me myself teaching.

OpenStudy (solomonzelman):

But teach you what ? \[application~~~of~~~~~\int\limits_{ }^{ } \int\limits_{ }^{ } \int\limits_{ }^{ }~ dx~dy~dz~~~~?\]

OpenStudy (yanasidlinskiy):

Omg. Are you for real? Look at your comment. \(\huge{\uparrow}\) "About"

OpenStudy (yanasidlinskiy):

Lol, Solomon.

OpenStudy (yanasidlinskiy):

ga;gb;abgjfgblagbraegbearbb. That's what "application of......" looks like to me:)xD

OpenStudy (solomonzelman):

That's a triple integral. I read a little about it in calc1 book.

OpenStudy (yanasidlinskiy):

Yea. That's what it looks like it:) Wanna teach me now?

OpenStudy (yanasidlinskiy):

*awkward*

OpenStudy (solomonzelman):

you want to learn integrals ?!

OpenStudy (solomonzelman):

\[\rm{Power~~rule:~}~~~~\int\limits_{ }^{ } x^n~~dx~~=~~\frac{x^{n+1}}{n}\]

OpenStudy (solomonzelman):

btw, the highest course I took is trigonometry.

OpenStudy (yanasidlinskiy):

Wow!:D Sure. Lets learn about that:)